OpenStudy (anonymous):

Opal deposited $2,925.90 into a savings account with an interest rate of 3.9% compounded twice a year. About how long will it take for the account to be worth $6,000?

OpenStudy (cwrw238):

use this formula ; 6000 = 2925.99( 1 + 0.039/2)^ 2t where t = number of years

OpenStudy (cwrw238):

thats 2925.90

OpenStudy (cwrw238):

do you goe to find value of t from that formula?

OpenStudy (anonymous):

I have to give the answer in years and months, I got 18 years I think, but I can't figure out the months.

OpenStudy (cwrw238):

sorry i got called away - i'll see what i get now

OpenStudy (anonymous):

Okay thank you!

OpenStudy (cwrw238):

18.593 years = 18 years 7 months as the interest is paid semi-annually your answer is 18 years 6 months

OpenStudy (cwrw238):

hmm maybe 18 yrs 7 months is a better answer as the value will be below 6000 after the payment at 18 years 6 months

OpenStudy (cwrw238):

ok ?

OpenStudy (anonymous):

The options are 18 years and ten months or 18 years and 8 months

OpenStudy (anonymous):

so 18 years and 8 months, right?

OpenStudy (cwrw238):

yes- its a bit over 7 so 18yr 8 months is right

OpenStudy (anonymous):

Okk, thank you! & I have another quick question !

OpenStudy (cwrw238):

works out o 18 yrs 7.2 months

OpenStudy (anonymous):

Edith purchased a stove for $950. It depreciates about 0.7% each year. What is the value of the stove after six years?

OpenStudy (cwrw238):

lol - i cant think of a formula for that use calculator and enter 950 - 0.07 * 950 = Ans - 0.07 * Ans = - same thing another 4 times sorry gotta go now

OpenStudy (cwrw238):

aww - heres the formula its .007 not o.07 Depreciated amount = 950(1 - 0.007)^6 = 910.79